If the concentration of a particular reactant is doubled and the rate of reaction is cut in half, what must be the order of the reaction with respect to that reagent? Enter your answer as an integer.
Answer #1

Answer: The rate of the reaction with respect to that particular reactant must be -1 (negative one)

Explanation:

Lets assume that initial concentration of a particular reactant is X and final concentration is 2X

lets say initial rate is R and final rate is R/2 and let the rate constant be K

(all these assumptions are according to the question)

then R = K x (X)n                                   ...(1)

and R/2 = K x (2X)n                               ...(2)

where n is the order of the reaction with respect to the particular reactant.

On dividing (1) by (2), we get

2 = (X)n/(2X)n = (1/2)n = 2-n

This further gives,

2 = 2-n

which says

-n = 1

n = -1